# Hessian Spectrum at the Ground State in Spherical Spin Glasses

### v1

@@ Abstract
We determine the bulk Hessian spectrum at approximate ground states in any spherical spin glass. Marginal stability, the presence of many near-zero Hessian eigenvalues, is shown to be equivalent to full replica symmetry breaking near overlap $1$ at zero temperature. This connection has long been implicit in the physics literature, which also links marginal stability to the performance of efficient algorithms. For even models, we prove the Hessian has no outlier eigenvalues, which has geometric consequences for low temperature Gibbs measures in the case that marginal stability is absent. Our proofs rely on interpolation bounds for certain vector spin glass models. For generic spin glasses, we give another more conceptual argument that full RSB near overlap $1$ implies marginal stability at low temperature.
@@ 1 Introduction
Spherical spin glass Hamiltonians are random disordered smooth functions in very high dimension. Their landscapes are understood to be rich and complicated, with exponentially many local maxima at a range of energy levels. We will be interested in the qualitative behavior around their near-global maxima. Does local uniform concavity hold near the extreme values, or could the Hessian be ill-conditioned? The former would imply that low temperature Gibbs measures are supported on isolated wells, and that the corresponding stationary dynamics remain trapped within them. By contrast an ill-conditioned marginally stable Hessian would allow for the possibility of a connected “manifold” of near maxima. In the physics literature, marginal stability at low-temperature is widely believed to be equivalent to full replica symmetry breaking (full RSB) near overlap $1$ at zero temperature, which is a property of the order parameter in the Parisi formula. We prove this equivalence for spherical spin glasses with even interactions, and derive consequences for Langevin dynamics and disorder chaos at low temperature whenever full RSB is absent.
@@
is approximately achieved. The limiting value of ${\mathrm{GS}}_{N}$ is given by Parisi’s formula. Following [CS17] (see also [CS92, Tal06a, Che13]), let ${\mathcal{N}}$ denote the set of non-decreasing, right-continuous functions $\zeta:[0,1)\to{\mathbb{R}}_{\geq 0}$ and let
@@ 1.1 Main Results
This result is stated as a proposition because a different formula for $\nabla H_{N}({\boldsymbol{\sigma}})$ follows using the ideas in [Sub24, Corollary 7] at zero temperature; see also [CPS23, CPS21] for Ising spin glasses. Our first step in proving (1.8) uses exactly this method, which yields a more abstract formula for $r(\xi)$ as the derivative of ${\mathrm{GS}}(\xi)$ along a perturbation of $\xi$. Applying the stationarity conditions for $(\zeta,L)$, we then deduce the more tractable (1.8). Regarding the bulk spectrum (1.9), known concentration estimates imply that uniformly for all ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$, the spectral measure of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$ is approximately a semi-circle density with radius $2\sqrt{\xi^{\prime\prime}(1)}$ and shifted by the rescaled radial derivative $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})=R({\boldsymbol{\sigma}},\nabla H_{N}({\boldsymbol{\sigma}}))$; see the proof of [Sub21, Lemma 3] (which is rephrased below as Lemma 1.11). Thus (1.9) follows routinely from (1.8).
@@
We pause here to emphasize two points. First, Corollary 1.5 relies crucially on the absence of Hessian outliers stated in (1.11): even a single near 00 outlier eigenvalue could result in $A_{\delta}$ having large components. Second, the full RSB assumption is essentially necessary: when $\lambda_{+}=0$, Corollary 1.5 evidently does not hold at any ${\boldsymbol{\sigma}}\in A_{\delta}$. Indeed, it is predicted in e.g. [CK94, Section 9] that under full RSB behavior, $\delta=O(1/N)$ may suffice for $A_{\delta}$ to have components of macroscopic diameter. Rigorously, it follows (e.g. by generic perturbations as in [Pan13b, Chapter 3.7]) that when $\xi$ is full RSB near overlap $1$ in the sense that $\zeta$ from (1.4) is strictly increasing on $[q,1]$, then for any $\delta>0$ and sufficiently large $N$, with high probability there exist ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\in A_{\delta}$ with $R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})=q$.
@@ 1.3 On Marginal Stability
See also [MW15, FU22]. It is separately understood, without clear mathematical justification, that “reasonable” optimization algorithms in high-dimension rapidly reach and then get stuck in the “manifold” of marginal states [CK94, Ken24]. As rigorous evidence for this, [HS24a] analyzed the optimal message-passing algorithms to optimize $H_{N}$ over ${\mathcal{S}}_{N}$ and showed the resulting outputs are marginally stable outside of the topologically trivial phase, even in the more general multi-species setting. Additionally, optimal stable algorithms to optimize mean-field spin glass Hamiltonians are now understood to be closely related to full RSB [Sub21, Mon21, AMS21, HS24b, AMS23b, HS23b, MZ24, JSS24]. Corollary 1.4 gives another rigorous link between full RSB and marginal stability, which however concerns low temperature statics rather than efficient algorithms.
@@ 1.4 Proof Ideas
The radial derivative formula (1.8) is proved using perturbative arguments in ${\mathbb{R}}^{N}$. The idea is to consider the maximum value of $H_{N}$ on a dilated spherical domain $t{\mathcal{S}}_{N}$. We take $|t-1|$ small (but independent of $N$), so the change in the maximum of $H_{N}$ is essentially given by the maximum or minimum radial derivative value at any approximate ground state (depending on the sign of $t-1$). On the other hand, it follows from (1.2) that the asymptotic ground state energy on a dilated sphere is given by ${\mathrm{GS}}(\xi_{t})$ for some slightly perturbed $\xi_{t}\approx\xi$. In this way, our proof of (1.8) first establishes an abstract formula for the radial derivative at any approximate ground state, expressed as a derivative of ${\mathrm{GS}}(\xi_{t})$. This relies on continuous differentiability properties of the spherical Parisi formula, which stem from the strict convexity of ${\mathcal{Q}}$. We then substantially simplify this formula using the stationarity conditions for the minimizing $(\zeta,L)$, yielding (1.8) and hence also (1.9).
The proofs of (1.11) and (1.12), which ensure the absence of Hessian outliers, are more delicate and given in Subsections 2.2 and 2.3 respectively. Here we perturb $H_{N}$ by considering augmented Hamiltonians of two or three replicas with constrained overlaps very close to $1$. By construction, the near-extrema for the augmented systems are obtained by starting from ${\boldsymbol{\sigma}}\in A_{\delta}$, and perturbing ${\boldsymbol{\sigma}}$ in opposite directions along a Hessian eigenvector. Thus in each case, precise estimates on the augmented ground state energy translate to bounds on the extreme eigenvalues at any approximate ground state. We find explicit interpolation parameters in the corresponding vector Parisi formulas, and perform a first-order Taylor expansion of the resulting bounds for the augmented ground state energy. This yields eigenvalue estimates which turn out to match the edge of the bulk spectrum from (1.9). Here $\xi$ is required to be even so the interpolation bounds for vector spin glasses are valid.
@@ 2.1 Radial Derivative at Approximate Ground States
We would like to show $t\mapsto{\mathrm{GS}}(\xi_{t})$ is continuously differentiable using the envelope theorem (see e.g. [DM24, Chapter 2.4]). The first step is to check the minimizer varies within a compact space; note that apriori the space ${\mathcal{K}}$ is not compact. Thus, let
@@
Lemma 2.1 shows ${\mathrm{GS}}(\xi_{s})$ is defined as an infimum over a fixed compact set ${\mathcal{K}}_{C}\subseteq{\mathcal{K}}$; Proposition 1.13 guarantees this minimizer is unique. This easily implies $s\mapsto(\zeta_{s},L_{t})$ is continuous. The first equality now follows from the envelope theorem [MS02, Theorem 2 and Corollary 4], assuming the latter derivative of $Q$ exists. Indeed dominated convergence shows it is given by the explicit formula above. ∎
@@
Next we apply Proposition 2.2 and standard concentration estimates to determine the radial derivative at any near-ground state, thus proving (1.8). The point is that, as explained at the start of this subsection, if $\xi$ is the mixture function for $H_{N}({\boldsymbol{\sigma}})$, then $\xi_{t}$ is the mixture function for $H_{N}(t{\boldsymbol{\sigma}})$. On the other hand, elementary calculus in ${\mathbb{R}}^{N}$ shows the radial derivative at ground states determines the change in the maximum value of $H_{N}$ from a small dilation of ${\mathcal{S}}_{N}$.
@@ 2.2 No Upward Outliers in Even Models
When $\xi$ is even, the limit ${\mathrm{GS}}_{2,\varepsilon}(\xi)=\operatorname*{p-lim}_{N\to\infty}{\mathrm{GS}}_{2,\varepsilon}(H_{N})$ exists and is given by a two-dimensional constrained generalization of the Parisi formula. Although we refrain from giving the notationally heavy general definitions, vector models of this type have played an important role in spin glass theory, being used in Talagrand’s original proof of the Parisi formula [Tal06b, Tal06a] and subsequent works [PT07, CHHS15, Che17, CHL18, AC18, CGPR19].
@@ Acknowledgements
Thanks to Brice Huang for helpful discussions, and for comments on a previous version of the paper. We also thank Andrea Montanari and Pierfrancesco Urbani for enlightening conversations.

### v2

# On Marginal Stability in Low Temperature Spherical Spin Glasses

@@ Abstract
We show marginal stability of near-ground states in spherical spin glasses is equivalent to full replica symmetry breaking at zero temperature near overlap $1$. This connection has long been implicit in the physics literature, which also links marginal stability to the performance of efficient algorithms. For even models, we prove the Hessian has no outlier eigenvalues, and obtain geometric consequences for low temperature Gibbs measures in the case that marginal stability is absent. Our proofs rely on interpolation bounds for vector spin glass models. For generic models, we give another more conceptual argument that full RSB near overlap $1$ implies marginal stability at low temperature.
@@ 1 Introduction
Spherical spin glass Hamiltonians are random disordered smooth functions in very high dimension. Their landscapes are understood to be rich and complicated, with exponentially many local maxima at a range of energy levels. We will be interested in the qualitative behavior around their near-global maxima. Does local uniform concavity hold near the extreme values, or could the Hessian be ill-conditioned? The former would imply that low temperature Gibbs measures are supported on isolated wells, and that the corresponding stationary dynamics remain trapped within them. By contrast an ill-conditioned marginally stable Hessian would allow for the possibility of a connected “manifold” of near maxima. In the physics literature, marginal stability at low-temperature is widely believed to be equivalent to full replica symmetry breaking (full RSB) near overlap $1$ at zero temperature, which is a property of the order parameter in the Parisi formula. We prove a strong form of this equivalence for spherical spin glasses with even interactions, and derive consequences for Langevin dynamics and disorder chaos whenever full RSB is absent.
@@
is approximately achieved. The limiting value of ${\mathrm{GS}}_{N}$ is given by Parisi’s formula. Following [CS17] (see also [CS92, Tal06a, Che13]), let ${\mathcal{N}}$ denote the set of non-decreasing, right-continuous functions $\zeta:[0,1)\to{\mathbb{R}}_{\geq 0}$ equipped with the vague topology, and define
@@ 1.1 Main Results
This result is stated as a proposition because a more abstract formula for $\nabla H_{N}({\boldsymbol{\sigma}})$, expressed as a derivative of ${\mathrm{GS}}(\xi)$, is essentially in [CS17, Remark 2] (or see [Sub24, Corollary 7]). Although we provide a detailed proof for completeness, the only novelty is an integration by parts using the stationarity conditions for $(\zeta,L)$, which yields the more tractable (1.8). Regarding the bulk spectrum (1.9), known concentration estimates imply that uniformly for all ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$, the spectral measure of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$ is approximately a semi-circle density with radius $2\sqrt{\xi^{\prime\prime}(1)}$ and shifted by the rescaled radial derivative $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})=R({\boldsymbol{\sigma}},\nabla H_{N}({\boldsymbol{\sigma}}))$; see the proof of [Sub21, Lemma 3] (which is rephrased below as Lemma 1.11). Thus (1.9) follows routinely from (1.8).
@@
We pause here to emphasize two points. First, Corollary 1.5 relies crucially on the absence of Hessian outliers stated in (1.11): even a single near-zero outlier eigenvalue could result in $A_{\delta}$ having large components. Second, the full RSB assumption is essentially necessary: when $\lambda_{+}=0$, Corollary 1.5 evidently does not hold at any ${\boldsymbol{\sigma}}\in A_{\delta}$. Indeed, it is predicted in e.g. [CK94, Section 9] that under full RSB behavior, $\delta=O(1/N)$ may suffice for $A_{\delta}$ to have components of macroscopic diameter. Rigorously, it follows (e.g. by generic perturbations as in [Pan13b, Chapter 3.7]) that when $\xi$ is full RSB near overlap $1$ in the sense that $\zeta$ from (1.4) is strictly increasing on $[q,1]$, then for any $\delta>0$ and sufficiently large $N$, with high probability there exist ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\in A_{\delta}$ with $R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})=q$.
@@ 1.3 On Marginal Stability
See also [MW15, FU22]. It is additionally believed that “reasonable” optimization algorithms in high-dimension rapidly reach and then get stuck in the “manifold” of marginal states [CK94, Ken24]; indeed [BAKZ22] recently conjectured that finding stable (i.e. non-marginal) local optima is computationally intractable in many disordered systems. As rigorous evidence for this, [HS24a] analyzed the optimal message-passing algorithms to optimize $H_{N}$ over ${\mathcal{S}}_{N}$ and showed the resulting outputs are marginally stable outside of the topologically trivial phase, even in the more general multi-species setting. Additionally, optimal stable algorithms to optimize mean-field spin glass Hamiltonians are now understood to be closely related to full RSB [Sub21, Mon21, AMS21, HS24b, AMS23b, HS23b, MZ24, JSS24]. Corollary 1.4 gives another rigorous link between full RSB and marginal stability, which however concerns low temperature statics rather than efficient algorithms.
@@ 1.4 Proof Ideas
The radial derivative formula (1.8) is proved using perturbative arguments in ${\mathbb{R}}^{N}$. The idea is to consider the maximum value of $H_{N}$ on a dilated spherical domain $t{\mathcal{S}}_{N}$. For $|t-1|$ small (but independent of $N$), the change in the maximum of $H_{N}$ is essentially given by the maximum or minimum radial derivative value at any approximate ground state (depending on the sign of $t-1$). On the other hand, it follows from (1.2) that the asymptotic ground state energy on a dilated sphere is ${\mathrm{GS}}(\xi_{t})$ for some slightly perturbed $\xi_{t}\approx\xi$. In this way, our proof of (1.8) first shows an abstract formula for the radial derivative at any approximate ground state; we then substantially simplify this formula using the stationarity conditions for the minimizing $(\zeta,L)$, yielding (1.8) and hence also (1.9).
The proofs of (1.11) and (1.12), which ensure the absence of Hessian outliers, are more delicate and given in Subsections 2.2 and 2.3 respectively. Here we perturb $H_{N}$ by considering augmented Hamiltonians of two or three replicas with constrained overlaps very close to $1$. By construction, the near-extrema for the augmented systems are obtained by starting from ${\boldsymbol{\sigma}}\in A_{\delta}$, and perturbing ${\boldsymbol{\sigma}}$ in opposite directions along a Hessian eigenvector. Thus in each case, precise estimates on the augmented ground state energy translate to bounds on the extreme eigenvalues at any approximate ground state. We find explicit interpolation parameters in the corresponding vector Parisi formulas, and perform a first-order Taylor expansion of the resulting bounds for the augmented ground state energy. This yields eigenvalue estimates which turn out to match the edge of the bulk spectrum from (1.9).
@@ 2.1 Radial Derivative at Approximate Ground States
We would like to show $t\mapsto{\mathrm{GS}}(\xi_{t})$ is in $C^{1}([1-\eta,1+\eta])$ using the envelope theorem. A general differentiation formula with respect to each coefficient $\gamma_{p}$ is stated in [CS17, Remark 2], analogously to the positive temperature case; this of course suggests a formula for $\frac{{\rm d}}{{\rm d}t}{\mathrm{GS}}(\xi_{t})$ by linearity. For completeness we give a careful proof, primarily checking that the minimizers remain within a suitable compact subset of ${\mathcal{K}}$ [^2] To illustrate the need for care, note that ${\mathcal{K}}$ is not compact, and $\zeta\mapsto\int_{0}^{1}\zeta(q){\rm d}q$ is lower semi-continuous but not continuous in the vague topology. Lower semi-continuity does not imply $\inf\limits_{|t-1|\leq\eta}\hat{\zeta}_{t}(1)>0$, even given that $t\mapsto(\zeta_{t},L_{t})\in{\mathcal{K}}$ is continuous and $\hat{\zeta}_{t}(1)>0$ for fixed $t$. Thus, let
@@
Lemma 2.1 shows ${\mathrm{GS}}(\xi_{s})$ is defined as an infimum over the compact set ${\mathcal{K}}_{C}\subseteq{\mathcal{K}}$; Proposition 1.13 ensures the minimizer is unique. This easily implies $s\mapsto(\zeta_{s},L_{t})$ is continuous. The first equality now follows from the envelope theorem [MS02, Theorem 2 and Corollary 4], assuming said derivative of ${\mathcal{Q}}$ exists. Indeed dominated convergence shows it is given by the explicit formula above. (As $C$ is fixed, there are no analytic issues for $q\approx 1$.) ∎
Next we show this formula coincides with $r(\xi)$ as defined in (1.5), using the stationarity conditions for $(\zeta,L)$.
@@
Next we apply Proposition 2.2 and standard concentration estimates to determine the radial derivative at any near-ground state, thus proving (1.8). The point is that, as explained at the start of this subsection, if $\xi$ is the mixture function for $H_{N}({\boldsymbol{\sigma}})$, then $\xi_{t}$ is the mixture function for $H_{N}(t{\boldsymbol{\sigma}})$. On the other hand, elementary calculus in ${\mathbb{R}}^{N}$ shows the radial derivative at ground states determines the change in the maximum value of $H_{N}$ from a small dilation of ${\mathcal{S}}_{N}$. (See [AC18a] for general related results.)
@@ 2.2 No Upward Outliers in Even Models
When $\xi$ is even, the limit ${\mathrm{GS}}_{2,\varepsilon}(\xi)=\operatorname*{p-lim}_{N\to\infty}{\mathrm{GS}}_{2,\varepsilon}(H_{N})$ exists and is given by a two-dimensional constrained generalization of the Parisi formula. Although we refrain from giving the notationally heavy general definitions, vector models of this type have played an important role in spin glass theory, being used in Talagrand’s original proof of the Parisi formula [Tal06b, Tal06a] and subsequent works [PT07, CHHS15, Che17, CHL18, AC18b, CGPR19].
@@ Acknowledgements
Thanks to Brice Huang for helpful discussions, and for comments on a previous version of the paper. We also thank Andrea Montanari, Pierfrancesco Urbani, and Lenka Zdeborová for enlightening conversations.